Let the joint pmf of X and Y be p(x, у) схуг, x-1,2,3, y-12. a) Find constant c that makes p(x, y) a valid joint pmf. c) Are X and Y independent? Justify d) Find P(X+Y> 3) and PCIX-YI # 1)
Answer #1

a)

for this to be valid: 3 sum_{y=1}^{2}p(x,y)=1

3sum_{y=1}^{2}p(x,y) =c(1*12+2*12+3*12+1*22+2*22+3*22)=c*30=1

c=1/30

b)

marginal pmf of X: f(x)=sum_{y=1}^{2}p(x,y) =(1/30)*(x*12+x*22)=5x/30 =x/6 for x=1,2,3

marginal pmf of y: f(y)=3 p(x, y) =(1/30)*(1*y2+2*y2+3*y2)=6y2/30 =y2/5 for y=1,2

c)

as P(x,y)=P(x)*P(y) ; therefore x and Y are independent

d)

P(X+Y>3)=P(X=2,Y=2)+P(X=3,Y=1)+P(X=3,Y=2)=(1/30)*(2*22+3*12+3*22)=23/30

P(|X-Y|eq1)

=P(X=1,Y=1)+P(X=2,Y=2)+P(X=3,Y=1)=(1/30)*(1*12+2*22+3*12)=12/30=2/5